# Barren-plateau free variational quantum simulation of $\mathbb{Z}_{2}$ lattice gauge theories

### v2

@@ Abstract
In this work, we design a variational quantum eigensolver (VQE) suitable for investigating ground states and static string breaking in a $\mathbb{Z}_{2}$ lattice gauge theory (LGT). We consider a two-leg ladder lattice coupled to Kogut-Susskind staggered fermions and verify the results of the VQE simulations using tensor network methods. We find that for varying Hamiltonian parameter regimes and in the presence of external charges, the VQE is able to arrive at the gauge-invariant ground state without explicitly enforcing gauge invariance through penalty terms. Additionally, experiments showing string breaking are performed on IBM’s quantum platform. Thus, VQEs are seen to be a promising tool for $\mathbb{Z}_{2}$ LGTs, and could pave the way for studies of other gauge groups. We find that the scaling of gradients with the number of qubits is favorable for avoiding barren plateaus. At the same time, it is not clear how to efficiently simulate the LGT using classical methods. Furthermore, strategies that avoid barren plateaus arise naturally as features of LGTs, such as choosing the initialization by setting the Gauss law sector and restricting the Hilbert space to the gauge-invariant subspace.
@@ I Introduction
Gauge theories, i.e. field theories with a local symmetry, are the theoretical foundation for the description of particle physics. At the same time, analytical access in the non-perturbative regime is often not possible. One way forward is lattice gauge theory (LGT), in which space-time is discretized on a lattice [1]. In this framework, matter fields are assigned to the lattice sites, while gauge fields are associated with the connecting links, all while maintaining gauge symmetry. In particular, this discretization allows for a numerical approach with sophisticated Monte Carlo methods.
While Monte Carlo methods have demonstrated excellent predictive power for static problems, they are limited by the sign problem in certain parameter regimes, leaving many relevant questions unanswered. This has triggered the ongoing effort to develop quantum and quantum-inspired methods for LGTs, such as tensor networks (TNs), to overcome these limitations. Current quantum processors are a long way from being used for realistic simulations of gauge theories. So far, successful demonstrations have been carried out for low-dimensional toy models [2, 3, 4, 5, 6, 7] of confinement [8, 9, 10], and scattering dynamics  [11, 12, 13, 14, 15]. For a review, see  [16, 17, 18].
In this work, we study a $\mathbb{Z}_{2}$ LGT [19] on a two-leg ladder geometry coupled to Kogut-Susskind fermions [20]. The two-leg ladder is an intermediate step between one and two dimensions. It is the simplest geometry that features magnetic plaquette interactions, and therefore all of the terms of the LGT Hamiltonian, while still being efficiently verifiable using tensor network techniques. The simplicity of the $\mathbb{Z}_{2}$ LGT is amenable to noisy intermediate-scale quantum (NISQ) devices, while still exhibiting complex phenomena, such as confinement.
Confinement is the mechanism that explains why isolated quarks or gluons are not observed in nature. The dynamical process of string breaking is fundamental to confining gauge theories, and is described by quantum chromodynamics. Using $\mathbb{Z}_{2}$ LGT as a toy model, we investigate static aspects of string breaking as a consequence of confinement. When two charges are confined in the $\mathbb{Z}_{2}$ LGT, by Gauss law they are necessarily connected by an electric flux tube. The flux tube experiences a linearly rising potential proportional to their distance. From a certain distance it becomes energetically favorable to break the flux string, and here this is done by forming an additional pair of charges.
@@
While we do not yet have the quantum hardware to answer this question, recent investigations of variational quantum algorithms have unearthed many issues. In particular, there seems to be a connection between the trainability of a VQE (the absence of barren plateaus) and efficient classical simulation [21]. Thus, it is important that proposals involving barren-plateau free VQEs consider the scaling with system size and the effects of noise.
In this work, we find features within the framework of $\mathbb{Z}_{2}$ LGT that are compatible with common strategies for avoiding barren plateaus, all the while being nontrivial to classically simulate. Warm starts are one such strategy, where one chooses to initialize close to a solution, where there may be narrow gorges [22]. Analogously, for investigations of LGT, it is necessary to initialize to some Gauss law sector, which thereby restricts the search landscape. The search process for a VQE is benefited by constraining to an explorable subspace of Hilbert space [23]; by gauge symmetry for LGTs, this is the subspace of the gauge-invariant, physical, states.
We draw upon these features using two types of ansatz circuits. One is restricted entirely to the symmetric subspace of physical states by being composed of gauge invariant terms. The other has overlap with states outside of the physical subspace, but nonetheless arrives at gauge invariant solutions over the course of the optimization. For the simulations we consider scaling with system size and numbers of shots, to find good convergence of ground states and we are able to witness static string breaking. Results are also obtained using IBM’s quantum hardware, where we perform experiments in the region of the transition into the string-broken phase. These findings underscore the usefulness of VQEs in the investigation of LGTs.
@@ II.1 Lattice gauge theories
Figure 1: Features of LGTs on the two-leg ladder lattice. The LGT specifies matter sites and links between them pertaining to the underlying gauge field. The arrows on the links show the orientation of the lattice. We have labeled $\hat{E}_{x}$ and $\hat{E}_{y}$ the orientations of the electric field and the ring labeled $\hat{P}$ indicates the magnetic plaquette term. For the vertex $\bm{l}$, we have indicated the notation for the adjacent link and matter site that is one unit vector $\hat{x}$ across. The red arrows show how we linearly arrange the qubits to perform the Jordan-Wigner transformation: we count sites from the leftmost lower corner, then traverse upwards, then rightwards along the lattice. The alternating colors of the matter sites illustrate the coupling to staggered fermions where the sign alternates. On this diagram we have also included an example string configuration, through the solid arrow connecting two charges that have been placed. The overlapping circles show the change in sign to indicating the placed charges.
The defining feature of gauge theories is invariance under local transformations. This leads to the presence of a local conservation law or a continuity equation that determines the interplay between the matter and gauge degrees of freedom. In the Hamiltonian formulation using temporal gauge, this results in an additional constraint known as Gauss law.
@@
The pure gauge part of the theory is constructed following Ref. [19] and the lattice notation for the two-dimensional ladder geometry we consider is illustrated in Fig. 1. When the underlying gauge group is $\mathbb{Z}_{2}$, the unitary operators acting on the links fulfill the $\mathbb{Z}_{2}$ algebra, so they can be re-written in terms of the Pauli matrices $X$ and $Z$. For the ladder geometry, the resulting pure gauge part of the Hamiltonian is given by:
@@
To couple the theory to fermionic matter, we consider Kogut-Susskind staggered fermions [20]. This formulation essentially corresponds to separating the spinor components to different lattice sites, resulting in alternating the signs of the masses and charges across the lattice as reflected in the orange and blue coloring in Fig. 1.
@@
The prefactor represents the different nature of the odd and even lattice sites, and we define the number operator $n_{l}=\phi_{l}^{\dagger}\phi_{l}$. Depending on the point of the lattice, the operators acting on the links are replaced by the boundary condition $X_{\bm{l},\bm{l}+\hat{x}}\!\rightarrow 1$.
@@
For simulations of LGTs, it is necessary for physical results to be gauge invariant, making it necessary to enforce Gauss law. One way to enforce gauge invariance is through penalty terms that are proportional to (3). Gauge-invariant states subject to gauge-invariant operators remain in this physical subspace, i.e., the symmetry remains intact. This subspace is still exponentially large, while being non-trivial to split into further gauge-invariant subspaces  [24, 25].
Incorporating terms proportional to (3) is also a means to introduce infinitely massive static charges to the theory. A consequence of Gauss law is that a pair of charges placed on matter sites in vicinity of each other are necessarily connected by a flux tube or string. These charges experience a linearly rising potential that is proportional to their separation.
From a certain distance onward, it becomes energetically favorable to break the string, which creates an additional pair of charges. This is known as static string breaking, and is depicted in Fig. 1, where a configuration of a string is shown between two sites with charges, indicated by the opposite coloring. Static string breaking and confinement has also been investigated for $\mathbb{Z}_{2}$ LGT in the following works [26, 27, 28, 10].
@@
To compute expectation values of this Hamiltonian on a quantum computer, we perform the Jordan-Wigner transformation [29] on the matter Hamiltonian (2) that couples to fermionic degrees of freedom. The gauge part and penalty terms already consist of Pauli operators. For this we must choose a linear ordering of the fermions. In Fig. 1, the ordering zig-zags from the bottom leg to the top, running across the rungs towards the left, as indicated by the red arrows. Subsequently, we translate the fermions to spins using $\phi_{\bm{l}}=\sideset{}{{}_{k<l}}{\prod}(iZ_{k})\sigma_{l}^{-}$, with $\sigma^{\pm}=\frac{1}{2}(X\pm iY)$ and $l$ being the linear index of the ordering of the fermionic fields as described.
@@ II.2 Variational quantum eigensolver
Circumstantially, there seems to be the additional issue that the trainability of a VQE implies that it is classically simulable. We can avoid barren plateaus by using ansatz circuits that explore smaller regions of Hilbert space related to the underlying problem structure. One way to reduce the dimensionality of the parameter space is to apply the symmetries of the system. Herein lies what enables efficient classical simulation – this reduction tends to identify some polynomially large subspace [21].
As such, it is possible to diagnose barren plateaus from the size of the dynamical Lie algebra (DLA) spanned by the gate generators. For the gates of the parametrized ansatz circuit, the DLA generates any of the arbitrary circuits that it can express. In [30] it is conjectured that the gradients of the cost function scale inverse-polynomially with the dimension of the DLA. However, while a small DLA may avoid barren plateaus, a polynomial-sized DLA implies the problem is classically simulable [31]. Classical simulation can be performed using techniques such as low-weight Pauli propagation [32, 33].
These are important considerations for ansatz design. A choice of ansatz that considers the available native gates, we call a hardware efficient ansatz (HEA). Whereas an ansatz based on unitaries formed of the problem Hamiltonian is known as the Hamiltonian variational ansatz (HVA). For the local Hamiltonian $H=H_{K}$, where each $H_{K}$ consists of commuting Pauli operators, the HVA can be written [34, 35] as:
@@
Initialization strategies to avoid barren plateaus have been proposed for both the HEA [36] and HVA [37]. However, such warm starts and initialization strategies show promise for VQEs, but are not guaranteed to find the global minimum, nor is there a general initialization strategy in subspaces containing the solution. This occurs for only a vanishing fraction of problem instances [22].
In [38] it is proven that local Hamiltonians training finite or logarithmic local-depth circuits (FLDC) will not encounter barren plateaus. Here the two-dimensional Toric code is considered, which is a variant of the pure gauge part of the $\mathbb{Z}_{2}$ LGT. In general FLDC cannot be classically simulated for two dimensions and higher.
@@ II.3 Tensor network states
MPS can be used as a family of variational ansätze to compute the ground state of a (local) Hamiltonian. To this end, the energy expectation value is variationally minimized by successively updating a tensor while keeping the others fixed. At each step, the optimal tensor is found by determining the ground state of an effective Hamiltonian describing the interaction of the site with its environment. Iterating this procedure until convergence, the final MPS provides an approximation for the ground state of the Hamiltonian. For our numerical calculations, we use the ITensor library where we run a standard two-site algorithm updating two neighboring tensors together [41, 42]. This allows for adaptively adjusting the bond dimension during the computation to obtain results with a certain precision. A detailed review on MPS methods can be found in Ref. [43].
@@ III.1 Choice of ansatz
One way to enforce and preserve gauge invariance is by constructing a gauge-preserving ansatz [44, 10], as shown in Fig. 2a. This ansatz, which we label GI, is a Hamiltonian variational ansatz, where the parametrized unitaries are each of the terms from the Hamiltonian Eq. (4). Upon transforming Eq. (2) to spin degrees of freedom using the Jordan-Wigner transformation acting on the relevant matter and link qubits, the kinetic term introduces the unitaries proportional to, e.g., $\exp{(-i\theta/2X\!\otimes\!Z\!\otimes\!Z\!\otimes\!Y)}$.
We also consider a simpler ansatz, that is presented in Fig. 2b and we refer to it as ZZ. The gates used in this ansatz are inspired by measurement-based quantum computing (MBQC) [45], consisting of single qubit rotation gates, and multi-qubit $Z$ rotations for the entangling layers. Such multi-qubit rotations can be implemented in MBQC by the introduction of one ancillary qubit [46]. The choice of simpler gates also makes this ansatz more hardware-efficient. The ansatz was designed by considering that the gauge part of the Hamiltonian in Eq. (1) has 4-body plaquette interactions and the kinetic part of the Hamiltonian in Eq. (2) introduces 3-body interactions along the edges connecting matter sites (see Fig. 1). We note that these are not truly 3-body terms as the transport law from gauge invariance involves intermediate matter qubits where the edge is not formed of adjacent qubits. These 3- and 4-body terms are accounted for through $R_{ZZZ}$ and $R_{ZZZZ}$ rotation gates on the relevant qubits. Furthermore this ansatz is gauge invariant when we initialize close to the angle $\theta\approx\pi$ for all the parametrized gates. Thus the ansatz has overlap with the gauge-invariant subspace and we can initialize the VQE in this space.
@@
We find good properties of the choice of ansätze. In Fig. 3 we show how the variance of the gradient of the cost function scales with system size and numbers of layers for the choice of ansatz circuit. For the simulations we have averaged over 100 samples and find favorable scaling of the variance up to system sizes of four plaquettes and here the Hamiltonian is the same as in Fig. 5. We also find that for both ansatz circuits, when scaling from one to two plaquettes, computing the DLA has too large a computational cost. The GI ansatz has similarities with the DLA of the matchgate circuit, which is known to be barren-plateau free and classically simulable [47]. However, the presence of $X$ gates make it so that the size of the DLA for the GI ansatz of one plaquette is $4080$, while the matchgate circuit on the same number of qubits ($8$) has dimension $120$. Another such example where the scaling of the variance of the cost function changes upon the inclusion of additional unitaries is shown in [30].
@@
In the noiseless case, low-weight Pauli propagation (PP) is a useful technique for computing expectation values where the problem is efficiently classically simulable [33]. However, for these simulations we find large errors and ineffective weight truncation. Importantly, these ansatz circuits use non-Clifford rotation gates, so that their entangling layers do not meet the criteria to use this method, but all the while evading barren-plateaus.
The LGT framework provides natural parallels with the features of VQEs that escape barren plateaus. For the average case it is not known how best to initialize in the cost landscape, while in LGT this is answered for by Gauss law sectors and the gauge invariant subspace. Gauss law also acts as a signal to when the simulation may be trapped in local minima, since Gauss law is violated for these states. For VQEs to work, the optimization uses symmetries of the problem, while here we have the gauge-invariant subspace. Furthermore, a problem is classically simulable if this subspace is polynomially sized, but we know the overall gauge-invariant subspace to be exponential in size [24, 25].
For the results, classical simulation of the VQE was performed using Pennylane [48], and to interact with IBM’s quantum platform, we have used Qiskit [49].

### v3

@@ Abstract
In this work, we design a variational quantum eigensolver (VQE) suitable for investigating ground states and static string breaking in a $\mathbb{Z}_{2}$ lattice gauge theory (LGT). We consider a two-leg ladder lattice with Kogut-Susskind staggered fermions and verify the results of the VQE simulations using tensor network methods. We find that for varying Hamiltonian parameter regimes and in the presence of external charges, the VQE is able to arrive at the gauge-invariant ground state without explicitly enforcing gauge invariance through penalty terms. Additionally, experiments showing string breaking are performed on IBM’s quantum platform. Thus, VQEs are seen to be a promising tool for $\mathbb{Z}_{2}$ LGTs, and could serve as a stepping stone toward studies of other gauge groups. We find that the scaling of gradients with the number of qubits is favorable for avoiding barren plateaus. Furthermore, strategies that avoid barren plateaus arise naturally as features of LGTs, such as choosing the initialization by setting the Gauss law sector and restricting the Hilbert space to the gauge-invariant subspace.
@@ I Introduction
Gauge theories, i.e. field theories with a local symmetry, are the theoretical foundation for the description of particle physics. At the same time, analytical access in the non-perturbative regime is often not possible. One way forward is lattice gauge theory (LGT), in which space-time is discretized on a lattice [Wilson1974]. In this framework, matter fields are assigned to the lattice sites, while gauge fields are associated with the connecting links, all while maintaining gauge symmetry. In particular, this discretization allows for a numerical approach with sophisticated Monte Carlo methods.
While Monte Carlo methods have demonstrated excellent predictive power for static problems, they are limited by the sign problem in certain parameter regimes, leaving many relevant questions unanswered. This has triggered the ongoing effort to develop quantum and quantum-inspired methods for LGTs, such as tensor networks (TNs), to overcome these limitations. Current quantum processors are a long way from being used for realistic simulations of gauge theories. So far, successful demonstrations have been carried out for low-dimensional toy models [Martinez2016, Atas2021, Atas2022, farrell2024scalable, farrell2024quantum, Klco2018, Klco2019, Kokail2018, meth2025simulating, Cochran2024, Cobos2025] of confinement [Mildenberger2025, Ciavarella2025, alexandrou2025realizingstringbreakingdynamics, Crippa2024a], and scattering dynamics  [Davoudi2024, Zemlevskiy2024, chai_fermionic_2025, chai_resource-efficient_2025]. For reviews, see, e.g. Refs. [Banuls2019, Banuls2020, Funcke2023a, Halimeh2025].
In this work, we study a $\mathbb{Z}_{2}$ LGT [Horn1979] on a two-leg ladder geometry coupled to Kogut-Susskind fermions [Kogut1975]. The two-leg ladder is an intermediate step between one and two dimensions. It is the simplest geometry that features magnetic plaquette interactions, and therefore all of the terms of the LGT Hamiltonian, while still being efficiently verifiable using tensor network techniques. The simplicity of the $\mathbb{Z}_{2}$ LGT is amenable to noisy intermediate-scale quantum (NISQ) devices, while still exhibiting complex phenomena, such as confinement.
Confinement is the mechanism that explains why isolated quarks or gluons are not observed in nature. The dynamical process of string breaking is fundamental to confining gauge theories, and is described by quantum chromodynamics. Using $\mathbb{Z}_{2}$ LGT as a toy model, we investigate static aspects of string breaking as a consequence of confinement. At large values of the coupling, when two charges are confined in the $\mathbb{Z}_{2}$ LGT, by Gauss law they are necessarily connected by an electric flux tube. Increasing their distance leads to a linearly rising potential. At large distances, it becomes energetically favorable to break the flux string, and here this is done by forming an additional pair of charges. In addition, at intermediate values of the coupling, a roughening transition can occur, which is a phase transition where the confining flux tube becomes rough, i.e., its transverse fluctuations become significant. This transition has recently been observed in TN simulations [marcantonio_roughening_2025].
@@
While we do not yet have the quantum hardware to answer this question, recent investigations of variational quantum algorithms have unearthed many issues. In particular, there seems to be a connection between the trainability of a VQE (the absence of barren plateaus) and efficient classical simulation [cerezo2024doesprovableabsencebarren]. Thus, it is important that proposals involving barren-plateau free VQEs consider the scaling with system size and the effects of noise.
In this work, we find features within the framework of $\mathbb{Z}_{2}$ LGT that are compatible with common strategies for avoiding barren plateaus, all the while being nontrivial to classically simulate. Warm starts are one such strategy, where one chooses to initialize close to a solution, where there may be narrow gorges [Puig2025] in the search landscape. Analogously, for investigations of LGT, it is necessary to initialize to some Gauss law sector, which thereby restricts the search landscape. The search process for a VQE is benefited by constraining to an explorable subspace of Hilbert space [Nakanishi19]; by gauge symmetry for LGTs, this is the subspace of the gauge-invariant, physical, states.
We draw upon these features using two types of ansatz circuits. One is restricted entirely to the symmetric subspace of physical states by being composed of gauge invariant terms. The other has overlap with states outside of the physical subspace, but nonetheless arrives at gauge invariant solutions over the course of the optimization. For the simulations we consider scaling with system size and numbers of shots, to find good convergence of ground states and we are able to witness static string breaking. Results are also obtained using IBM’s quantum hardware, where we perform experiments in the region of the transition into the string-broken phase. These findings underscore the usefulness and suitability of VQEs in the investigation of LGTs.
@@ II.1 Lattice gauge theories
Figure 1: Features of LGTs on the two-leg ladder lattice. The LGT specifies matter sites and links between them pertaining to the underlying gauge field. The arrows on the links show the orientation of the lattice. We have labeled $\hat{E}_{x}$ and $\hat{E}_{y}$ the orientations of the electric field and the ring labeled $\hat{P}$ indicates the magnetic plaquette term. For the vertex $\bm{l}$, we have indicated the notation for the adjacent link and matter site that is one unit vector $\hat{x}$ across. The red arrows show how we linearly arrange the qubits to perform the Jordan-Wigner transformation: we count sites from the leftmost lower corner, then traverse upwards, then rightwards along the lattice. The alternating colors of the matter sites illustrate the coupling to staggered fermions where the sign alternates. On this diagram we have also included an example string configuration, through the solid arrow connecting two charges that have been placed. The overlapping circles show the change in sign, indicating the placed charges.
The defining feature of gauge theories is invariance under local transformations. This leads to the presence of a local conservation law or a continuity equation that determines the interplay between the matter and gauge degrees of freedom. In the Hamiltonian formulation using temporal gauge, this results in an additional constraint known as Gauss law [Kogut1975].
@@
The pure gauge part of the theory is constructed following Ref. [Horn1979] and the lattice notation for the two-dimensional ladder geometry we consider is illustrated in Fig. 1. When the underlying gauge group is $\mathbb{Z}_{2}$, the unitary operators acting on the links fulfill the $\mathbb{Z}_{2}$ algebra, so they can be re-written in terms of the Pauli matrices $X$ and $Z$. For the ladder geometry, the resulting pure gauge part of the Hamiltonian is given by:
@@
To couple the theory to fermionic matter, we consider Kogut-Susskind staggered fermions [Kogut1975]. This formulation essentially corresponds to separating the spinor components to different lattice sites, resulting in alternating the signs of the masses and charges across the lattice as reflected in the orange and blue coloring in Fig. 1.
@@
The prefactor represents the different nature of the odd and even lattice sites, and we define the number operator $n_{l}=\phi_{l}^{\dagger}\phi_{l}$. Following Ref. [alexandrou2025realizingstringbreakingdynamics], we interpret $q_{\bm{l}}=1$ as the absence of a charge, whereas a value of 1-1 corresponds to a static charge at site $\bm{l}$. Depending on the point of the lattice, the operators acting on the links are replaced by the boundary condition $X_{\bm{l},\bm{l}+\hat{x}}\!\rightarrow 1$ (see Ref. [alexandrou2025realizingstringbreakingdynamics] for details).
@@
For simulations of LGTs, it is necessary for physical results to be gauge invariant, making it necessary to enforce Gauss law. One way to enforce gauge invariance is through penalty terms that are proportional to (3). Gauge-invariant states subject to gauge-invariant operators remain in this physical subspace, i.e., the symmetry remains intact. This subspace is still exponentially large, while being non-trivial to split into further gauge-invariant subspaces  [Buividovich_2009, Tagliacozzo2014].
Incorporating terms proportional to (3) allows for enforcing Gauss law, where infinitely massive static charges can be introduced to the theory by setting the corresponding $q_{\bm{l}}=-1$. A consequence of Gauss law is that a pair of charges placed on matter sites in vicinity of each other are necessarily connected by a flux tube or string. These charges experience a linearly rising potential that is proportional to their separation.
From a certain distance onward, it becomes energetically favorable to break the string, which creates an additional pair of charges. This is known as static string breaking, and is depicted in Fig. 1, where a configuration of a string is shown between two sites with charges, indicated by the opposite coloring. Static string breaking and confinement has also been investigated for $\mathbb{Z}_{2}$ LGT in the following works [Lumia2022Z2, Xu2025, marcantonio_roughening_2025, alexandrou2025realizingstringbreakingdynamics].
@@
To compute expectation values of this Hamiltonian on a quantum computer, we perform the Jordan-Wigner transformation [Hamer1997Series] on the matter Hamiltonian (2) that couples to fermionic degrees of freedom. For this we must choose a linear ordering of the fermions. The gauge part and penalty terms already consist of Pauli operators. In Fig. 1, the ordering zig-zags from the bottom leg to the top, running across the rungs towards the left, as indicated by the red arrows. Subsequently, we translate the fermions to spins using $\phi_{\bm{l}}=\sideset{}{{}_{k<l}}{\prod}(iZ_{k})\sigma_{l}^{-}$, with $\sigma^{\pm}=\frac{1}{2}(X\pm iY)$ and $l$ being the linear index of the ordering of the fermionic fields as described.
@@ II.2 Variational quantum eigensolver
Circumstantially, there seems to be the additional issue that the trainability of a VQE implies that it is classically simulable. We can avoid barren plateaus by using ansatz circuits that explore smaller regions of Hilbert space related to the underlying problem structure. One way to reduce the dimensionality of the parameter space is to apply the symmetries of the system. Herein lies what enables efficient classical simulation – this reduction tends to identify some polynomially large subspace [cerezo2024doesprovableabsencebarren].
As such, it is possible to diagnose barren plateaus from the size of the dynamical Lie algebra (DLA) spanned by the gate generators. For the gates of the parametrized ansatz circuit, the DLA generates any of the arbitrary circuits that it can express. In [Larocca2022diagnosingbarren] it is conjectured that the variance of gradients of the cost function scale inverse-polynomially with the dimension of the DLA. However, while a small DLA may avoid barren plateaus, a polynomial-sized DLA implies the problem is classically simulable [goh2025liealgebraicclassicalsimulationsquantum]. Classical simulation can be performed using the DLA [goh2025liealgebraicclassicalsimulationsquantum], or techniques such as low-weight Pauli propagation [schuster2024polynomialtimeclassicalalgorithmnoisy, angrisani2024classicallyestimatingobservablesnoiseless].
These are important considerations for ansatz design. A choice of ansatz that considers the available native gates, we call a hardware efficient ansatz (HEA). Whereas an ansatz based on unitaries formed of the problem Hamiltonian is known as the Hamiltonian variational ansatz (HVA). For the local Hamiltonian $H=H_{K}$, where each $H_{K}$ consists of commuting Pauli operators, the HVA can be written [Wecker2015, Wiersema2020] as:
@@
Initialization strategies to avoid barren plateaus have been proposed for both the HEA [Park:2024rim] and HVA [Park2024hamiltonian]. However, while warm starts and related initialization strategies show promise for VQEs, they are not guaranteed to find the global minimum. Nor does there exist a general initialization strategy in subspaces containing the solution. This occurs for only a vanishing fraction of problem instances [Puig2025].
In [Zhang2024Absence] it is proven that local Hamiltonians training finite or logarithmic local-depth circuits (FLDC) will not encounter barren plateaus. Here the two-dimensional Toric code is considered, which is a variant of the pure gauge part of the $\mathbb{Z}_{2}$ LGT. In general FLDC cannot be classically simulated for two dimensions and higher.
@@ II.3 Tensor network states
MPS can be used as a family of variational ansätze to compute the ground state of a (local) Hamiltonian. To this end, the energy expectation value is variationally minimized by successively updating a tensor while keeping the others fixed. At each step, the optimal tensor is found by determining the ground state of an effective Hamiltonian describing the interaction of the site with its environment. Iterating this procedure until convergence, the final MPS provides an approximation for the ground state of the Hamiltonian. For our numerical calculations, we use the ITensor library where we run a standard two-site algorithm updating two neighboring tensors together [Fishman_2022, Fishman2022b]. This allows for adaptively adjusting the bond dimension during the computation to obtain results with a certain precision. A detailed review on MPS methods can be found in Ref. [Schollwoeck2011].
@@ III.1 Choice of ansatz
One way to enforce and preserve gauge invariance is by constructing a gauge-preserving ansatz [mazzola2021gauge, alexandrou2025realizingstringbreakingdynamics], as shown in Fig. 2a. This ansatz, which we label GI, is a Hamiltonian variational ansatz, where the parametrized unitaries are each of the terms from the Hamiltonian Eq. (4). Upon transforming Eq. (2) to spin degrees of freedom using the Jordan-Wigner transformation acting on the relevant matter and link qubits, the kinetic term introduces the unitaries proportional to, e.g., $\exp{(-i\theta/2X\!\otimes\!Z\!\otimes\!Z\!\otimes\!Y)}$.
We also consider a simpler ansatz, that is presented in Fig. 2b and we refer to it as ZZ. The gates used in this ansatz are inspired by measurement-based quantum computing (MBQC) [mbqc_def], consisting of single qubit rotation gates, and multi-qubit $Z$ rotations for the entangling layers. Such multi-qubit rotations can be implemented in MBQC by the introduction of one ancillary qubit [qin2024applicability]. The choice of simpler gates also makes this ansatz more hardware-efficient. The ansatz was designed by considering that the gauge part of the Hamiltonian in Eq. (1) has 4-body plaquette interactions and the kinetic part of the Hamiltonian in Eq. (2) introduces 3-body interactions along the edges connecting matter sites (see Fig. 1). We note that these are not truly 3-body terms as the transport law from gauge invariance involves intermediate matter qubits where the edge is not formed of adjacent qubits. These 3- and 4-body terms are accounted for through $R_{ZZZ}$ and $R_{ZZZZ}$ rotation gates on the relevant qubits. Furthermore this ansatz is gauge invariant when we initialize close to the angle $\theta\approx\pi$ for all the parametrized gates. Thus the ansatz has overlap with the gauge-invariant subspace and we can initialize the VQE in this space.
@@
We find good properties of the choice of ansätze. In Fig. 3 we show how the variance of the gradient of the cost function scales with system size and numbers of layers for the choice of ansatz circuit. For the simulations we have averaged over 100 samples and find favorable scaling of the variance up to system sizes of four plaquettes and here the Hamiltonian is the same as in Fig. 6.
For both ansatz circuits, we find that when scaling from one to two plaquettes, computing the DLA has too large a computational cost. The GI ansatz has similarities with the DLA of the matchgate circuit, which is known to be barren-plateau free and classically simulable [diaz2023showcasingbarrenplateautheory]. However, the presence of $X$ gates make it so that the size of the DLA for the GI ansatz of one plaquette is $4080$, while the matchgate circuit on the same number of qubits ($8$) has dimension $120$. Another example where the scaling of the variance of the cost function changes upon the inclusion of additional unitaries is shown in [Larocca2022diagnosingbarren].
In order to consider some scaling relation of the DLA, we consider the non-physical scenario of adding fermions to the two-dimensional system one-by-one. This amounts to beginning with the smallest LGT, where we have two fermion qubits connected by a link qubit and the system size is increased by adding a fermion qubit alongside a respective link qubit. Up to $7$ qubits, the two-leg ladder is the same as the one-dimensional chain, so we cannot use these points to compute the scaling. Results for the dimension of the DLA, $\dim\mathfrak{g}$ and scaling is shown in Fig. 5. Here we see that the one-dimensional problem of both GI and ZZ ansatz show exponential scaling, and show some data for the two-leg ladder lattice.
Figure 5: Scaling of the dimension of the DLA. For increasing number of qubits of the GI and ZZ ansatz and the one- and two-dimensional LGT. Here we include an exponential and polynomial fit for the one-dimensional cases. For the GI ansatz the exponential fitting is given by $\dim(\mathfrak{g})\approx$3.49\text{\times}{10}^{-1}$2^{1.50n}$, where $n$ is the number of qubits, and the mean squared error (MSE) $10.56$, while the polynomial fit is $\dim(\mathfrak{g})\approx$5.71\text{\times}{10}^{-7}$n^{10.33}$ with MSE $8550.98$. Whereas for the ZZ ansatz the exponential fit is $\dim(\mathfrak{g})\approx$5.47\text{\times}{10}^{-1}$2^{1.98n}$ with MSE $1.63$, and then the polynomial fit is $\dim(\mathfrak{g})\approx$1.04\text{\times}{10}^{-3}$n^{8.17}$ with MSE $258.65$. Up to $7$ qubits, the one- and two-dimensional cases are the same, as the 8th qubit introduces a plaquette term. With this additional term, the DLA is larger than for the one-dimensional case.
@@
The LGT framework provides natural parallels with the features of VQEs that escape barren plateaus. For the average case it is not known how best to initialize in the cost landscape, while in LGT this is answered for by Gauss law sectors and the gauge invariant subspace. Gauss law also acts as a signal to when the simulation may be trapped in local minima, since Gauss law is violated for these states. For VQEs to work, the optimization leverages symmetries of the problem, and here this is the gauge-invariant subspace. Furthermore, a problem is classically simulable if this subspace is polynomially sized, but we know the overall gauge-invariant subspace to be exponential in size [Buividovich_2009, Tagliacozzo2014].
For the results, classical simulation of the VQE was performed using Pennylane [bergholm2022pennylaneautomaticdifferentiationhybrid], and to interact with IBM’s quantum platform, we have used Qiskit [javadiabhari2024quantumcomputingqiskit].